# AI-enabled tuberculosis screening in a high-burden setting using cough sound analysis and speech foundation models

### v1

@@ Abstract
Background Artificial intelligence (AI) systems can detect disease-related acoustic patterns in cough sounds, offering a scalable and cost-effective approach to tuberculosis (TB) screening, especially in high-burden, low-resource settings with limited access to care. However, prior studies have been constrained by small datasets, under-representation of symptomatic non-TB patients, reliance on simple machine learning models, and recordings collected under idealised conditions.
Methods We enrolled 512 participants at two hospitals in Zambia, categorised into three groups: bacteriologically confirmed TB (TB+), symptomatic patients with non-TB respiratory diseases (OR), and healthy controls (HC). Usable cough recordings, demographic, and clinical data were obtained from 500 participants. Deep learning classifiers based on speech foundation models were trained on cough sound recordings to predict diagnostic categories. The best-performing classifier, trained on 3-second cough segments, was further evaluated in combination with demographic and clinical data.
Findings The best-performing audio-only classifier achieved an AUROC of 85.2% for distinguishing TB+ coughs from all other participants (TB+/ Rest) and 80.1% for TB+ versus symptomatic OR participants (TB+/ OR). Combining demographic and clinical features improved performance, with AUROC reaching 92.1% for TB+/ Rest and 84.2% for TB+/ OR. At a probability threshold of 0.38, the sensitivity and specificity of the multimodal model were 90.3% and 73.1% for TB+/ Rest, and 80.6% and 73.1% for TB+/ OR.
Interpretation Cough sound analysis using speech foundation models, particularly when combined with demographic and clinical data, demonstrated strong potential as a TB triage tool and met the WHO target product profile benchmarks. The model was robust to potential confounding factors including background noise, recording time, and device variability, as shown through adversarial testing and stratified analyses. These findings support the model’s ability to capture disease-related acoustic patterns rather than artefactual cues. Further external validation across diverse regions and case definitions, including subclinical TB, is essential before clinical deployment.
Funding UK Higher Education Innovation Fund and UK Engineering and Physical Sciences Research Council (EPSRC) Impact Acceleration Account.
Copyright © 2025 The Author(s). Published by Elsevier Ltd. This is an Open Access article under the CC BY 4.0 license.
journal: Lancet Digital Health
Introduction
@@
Tuberculosis (TB) remains one of the most pressing global health challenges, with an estimated 10.6 million new cases and 1.3 million deaths reported in 2023 [1]. Despite being both preventable and curable, TB continues to be the leading cause of death from a single infectious agent. A significant proportion of TB-related deaths stems from undiagnosed and untreated cases – approximately 25% of all estimated TB cases in 2023 [1]. Systematic screening is essential to close the TB case detection gap, which facilitates earlier diagnosis, reduces community transmission, and improves treatment outcomes [2]. However, current World Health Organisation (WHO)-approved screening tools, including symptoms screening, chest X-ray (CXR), C-reactive protein (CRP), and molecular WHO-recommended rapid diagnostic (mWRD), have limitations that restrict their effectiveness and scalability. Symptom-based screening offers only moderate sensitivity (71%) and specificity (64%) [3], while CXR requires costly infrastructure, with limited availability in many low-resource settings. CRP testing shows diagnostic utility primarily among people living with HIV (PLHIV), and mWRD, although valuable, are constrained by moderate sensitivity (69%). It is only recommended for adolescents and adults living with HIV and requires a robust sample transportation network [4, 5, 6].
Research in context
Evidence before this study
We searched PubMed on 22nd May 2025 using the terms: (“cough*”) AND (“artificial intelligence*” OR “automatic*” OR “classifier*” OR “machine learning”) AND (“TB” OR “Tuberculosis” [MeSH]). Of the 39 results, only studies that evaluated the performance of AI enabled cough sound analysis against a TB microbiological reference standard were considered. Studies whose corresponding publications were not peer reviewed were excluded. Four studies that met the above criteria were identified. All four studies reported on the area under the receiver operating characteristic (ROC) curve (AUROC), sensitivity and specificity. Performance of the models was high across the studies. However, the risk of bias was high as two of the studies had very small numbers of participants, data biases due to the limited inclusion of symptomatic non-TB patients, the use of simplistic machine learning models, and data collection in idealised, quiet conditions.
Added value of this study
We developed an AI enabled cough analysis model for TB screening using a large dataset of cough recordings from 500 participants across three categories: bacteriologically confirmed TB patients, patients with respiratory diseases other than TB, and healthy controls. Our best performing model demonstrated a high performance, achieving an AUROC of 92.1%, sensitivity of 90.3% and specificity of 73.1% when demographic and clinical parameters were incorporated. We also demonstrate that the model’s classification relies on the cough sound itself as opposed to exploiting characteristics of the noise background. When the model was tested using background noise, the AUROC was 58.6%. Finally, we report differences in model performance when using cough recordings collected using high-quality microphones and those collected via mobile phone, highlighting practical considerations for real-world deployment.
Implications of all the available evidence
AI-enabled cough analysis has potential as a low-cost TB triage or self-screening tool. Rigorous external validation in community and clinical settings is needed to ensure generalisability, reliability, and safety before integration into screening programmes.
@@
Advances in artificial intelligence (AI) offer opportunities to improve TB screening and diagnosis by enhancing accuracy, efficiency, and accessibility across various healthcare domains, including medical imaging [7, 8, 9], disease detection [10, 11, 12, 13], and sleep-disordered breathing monitoring [14]. AI-powered CXR interpretation tools were recommended by WHO in 2021 as alternatives to human readers for TB screening and triage [2]. However, their use remains constrained by the limited availability of CXR equipment [5, 6]. These challenges highlight the need for alternative AI applications in TB screening and triage [2], particularly those that exploit other diagnostic substrates such as audio signals and wearable sensor data, which may be more accessible and scalable in low-resource settings.
@@
One promising alternative is AI analysis of cough sounds. Chronic cough, a dominant symptom of TB, is caused by inflammation of the airways. It is hypothesised that TB-related coughs carry distinct acoustic features that can be differentiated from coughs due to other causes [15]. Acoustic AI models offer the potential to objectively identify these features, thus reducing inter-listener variability and improving diagnostic consistency [16]. Building on this hypothesis, recent studies have explored AI-enabled cough sound analysis as a screening tool for TB [15, 17, 18, 19]. This approach aligns with clinical practice, where clinicians already use cough characteristics such as tone, pitch, intensity, and duration of coughs, alongside associated symptoms, to guide diagnosis [20, 21]. However, prior research in this area has been constrained by several limitations, including small sample sizes [15, 17], data biases due to inadequate inclusion of symptomatic individuals without TB [15, 17, 18], the use of relatively simple machine learning models [15, 22], and data collected in controlled, quiet environments that do not reflect real-world conditions [15, 19].
@@
In this study, we developed and evaluated an AI-enabled cough sound analysis model specifically designed to overcome these limitations. We hypothesised that, with a sufficiently large and well-balanced dataset, machine learning models could learn to recognise spectral and temporal features in cough sounds that reliably distinguish TB from non-TB cases, even when recorded using ambient sound conditions and widely available hardware. Our objective was to assess model performance in a high-burden setting, thereby advancing the development of practical and scalable tools for TB screening. Such tools could be particularly valuable in resource-limited health systems and in high-throughput contexts such as immigration screening centres.
Methods
Study design and participants
@@ Methods
We conducted a cross-sectional diagnostic accuracy study in accordance with the Standards for Reporting of Diagnostic Accuracy Studies (STARD) guidelines [23]. Ethical approval was obtained from the University of Zambia Biomedical Research Ethics Committee (approval number 3648-2023). Written informed consent was obtained from all participants prior to enrolment.
@@
A multi-stage classification pipeline (Figure 1) was developed using state-of-the-art pre-trained speech foundation models. These models, including Wav2Vec2 [25], WavLM [26], HuBERT [27], Data2Vec [28], and Whisper [29], accept raw audio waveform as input and generate contextualised audio embeddings that capture complex acoustic patterns. To adapt each model for three-way classification (TB+, OR, HC) in this study, we appended a classifier head comprising a global average pooling layer, dropout (rate = 0.5), and a softmax output layer. Fine-tuning was conducted using stratified 10-fold cross-validation with the following parameters: 5 training epochs, batch size = 8, learning rate = 3e-5, warm-up ratio = 0.1, and gradient accumulation over 8 steps.
@@
Role of the funding source
This study was funded by UK Higher Education Innovation Fund and UK Engineering and Physical Sciences Research Council (EPSRC) Impact Acceleration Account. The funders of the study had no role in study design, data collection, data analysis, data interpretation, or writing of the report.
Results
[…]
@@ Results
Table 2: Comparison of AUROC with different classifier architectures and duration of the audio signal input. Best results are in bold.
@@
TB+/ Rest Audio Alone 81.5% 75.4% 69.7% 73.2% 72.1% 74.1%
(HIV+) Audio + All Info 91.8% 81.9% 90.9% 94.7% 71.5% 87.8%
Using the Wav2Vec2-based classifier and 3-second audio input, we evaluated classification performance when demographic or clinical features (gender, age, BMI, or symptom presence) are available (Table 3). The ROC curves in different tasks are shown in Figure 2. With audio alone, classification performance was consistently better for the TB+/ HC task than for TB+/ OR. Specificity was similar across both tasks (73.1% vs. 73.2%), but sensitivity (92.6% vs. 72.0%), PPV (71.8% vs. 66.7%), NPV (93.1% vs. 77.8%) and F1-score (80.9% vs. 69.1%) were all notably higher for TB+/ HC. These differences suggest greater acoustic overlap between TB and other symptomatic respiratory conditions than between TB and healthy controls. When the OR and HC groups are combined in the TB+/ Rest task, sensitivity is 82.3%, specificity is 73.2%, PPV is 82.0%, NPV is 73.5% and F1-score is 82.1%. Adding individual demographic or clinical features to audio inputs resulted in modest improvements across all tasks. The most substantial gains were observed when all features were combined. For TB+/ Rest, the inclusion of all supplemental information increased AUROC from 85.2% to 92.1%, sensitivity from 82.3% to 85.6%, and specificity from 73.2% to 83.1%. Similarly, F1-score improved from 82.1% to 86.9%. In the more challenging TB+/ OR task, where both groups were symptomatic, AUROC increased from 80.1% to 84.2%, and specificity from 73.2% to 83.1%, although sensitivity declined slightly from 72.0% to 71.3%. Overall, combining audio with demographic and clinical data led to consistent performance improvements across most metrics, with the strongest effects seen in the TB+/ Rest task.
@@
Table 4 shows model performance with different classification thresholds ranging from 0.36 to 0.50. For the TB+/ Rest task, a threshold of 0.38 offered a strong balance of sensitivity (90.3%) and specificity (73.1%), meeting the WHO Target Product Profile for TB triage tests, which recommends >90% sensitivity and >70% specificity for screening tools. Higher thresholds improved specificity (up to 83.1% at 0.50) at the cost of reduced sensitivity (85.6% at 0.50). For TB+/ OR, performance was generally lower than TB+/ Rest. At the 0.38 threshold, sensitivity was 80.6% and specificity was 73.1%. As the threshold increased, specificity improved (up to 83.1%) while sensitivity dropped (71.3% at 0.50). In the TB+/ HC task, the model achieved perfect sensitivity (100%) across all thresholds tested, with specificity increasing from 70.2% to 83.1% between thresholds 0.36 and 0.50.
Table 4: Comparison of results for the proposed system (audio + all demographic information) with different threshold scores. Bracketed figures show 95% confidence intervals based on 10,000 repeats. The results meeting the WHO Target Product Profile for TB triage tests are in bold.
@@
Table 5: Effect of different noise training and testing conditions using the Wav2Vec2-based audio classifier.
@@
Non-Cough Background Non-Cough Background 69.9%
Cough Sound Non-Cough Background 58.6%
@@
When both trained and tested using white noise as audio input, the model performed at around the chance level (AUROC = 56.2%), indicating no meaningful discriminative capacity in the absence of cough sound (Table 5). In contrast, training and testing on non-cough acoustic background led to a significant increase in performance (AUROC = 69.9%), which suggests the presence of some background noise patterns associated with different participant groups. However, this performance remained well below that observed when trained and tested on cough sounds (AUROC = 85.2%), confirming that cough acoustics carry the primary discriminatory information. Notably, when the model was trained on cough sounds and tested on background noise, AUROC values again approached chance (58.6%), further underscoring the specificity of the learned representations to cough content rather than background artefacts.
@@
The spectral energy distribution of cough segments and background noise across groups can be visualised in the long-term average spectrum (LTAS) (Figure 3). Cough LTAS profiles (left panel) were broadly similar across TB+, OR, and HC groups, with only minor differences: TB+ exhibiting slightly lower energy around 200 Hz and higher energy above 1.5 kHz. This indicates that group-level differences in cough acoustics are not prominently expressed in the long-term spectral domain, suggesting that temporal features may offer greater diagnostic value. In contrast, background noise LTAS (right panel) showed clearer group differences, with TB+ recordings exhibiting a flatter, higher-energy spectrum across frequencies, likely due to contextual (e.g., recording environment or timing) rather than physiological factors. However, adversarial testing (Table 5) showed that any such confounding is minimal, as near-chance performance was reported when models were trained on cough sounds and tested on background noise.
@@
Table 6: Performance of Wav2Vec2 classifiers (with and without additional demographic information) using mobile phone test data.
@@
Discussion
This study demonstrates the potential of AI-enabled cough sound analysis using speech foundation models to identify TB from short cough recordings. Our multimodal model achieved performance that meets the WHO target product profile benchmarks for TB triage tools in distinguishing TB-positive cases from individuals without TB (TB+/Rest, AUROC 92.1%, sensitivity 90.3%, specificity 73.1%). Based on a well-balanced cohort of 500 participants (201 TB+, 150 symptomatic non-TB, 149 healthy controls), these findings highlight the promise of cough-based AI as a scalable, non-invasive, and low-cost screening tool in high-burden settings with limited diagnostic infrastructure.
Compared with previous studies, our model achieved robust and reproducible performance despite real-world recording conditions. Direct comparisons are limited due to key methodological differences across studies. Prior research in this area has faced several constrains, including smaller or imbalanced datasets [15, 17], limited inclusion of symptomatic non-TB patients [15, 17, 18], the use of relatively simple machine learning models [15, 22], and data collection in controlled, quiet environments [15, 19], leading to potential bias and inflated accuracy.
Our findings align with real-world clinical practice and prior studies [31], demonstrating that the inclusion of demographic and clinical parameters, such as age, gender, BMI, and symptom presence, can enhance model performance. For the TB+/ Rest task, model sensitivity improved substantially, primarily driven by the inclusion of asymptomatic healthy controls. As this subgroup lacks symptoms, the distinction between TB and non-TB cases becomes clearer, enabling the model to more accurately detect TB-positive individuals. In contrast, for the TB+/ OR task, where both groups were symptomatic, clinical features contributed primarily to improved specificity, allowing the model to better distinguish TB from other respiratory illnesses.
The model’s lower performance in the TB+/ OR task compared to TB+/ HC highlights the clinical complexity of differentiating TB from other respiratory conditions. Including symptomatic non-TB participants is crucial for realistic rigorous performance assessment. Prior studies have often overlooked this group, or employed a small imbalanced dataset, which could lead to inflated accuracy estimates [15, 17, 18]. For example, Sharma et al. [19] reported highly variable classifier performance (AUROC 61–86%), largely due to training on imbalanced datasets (103 TB patients and 46 non-TB participants).

### v2

# Deep Learning for Tuberculosis Screening in a High-burden Setting using Cough Analysis and Speech Foundation Models

@@ Abstract
Artificial intelligence (AI) systems can detect disease-related acoustic patterns in cough sounds, offering a scalable and cost-effective approach to tuberculosis (TB) screening in high-burden, resource-limited settings. Previous studies have been limited by small datasets, under-representation of symptomatic non-TB patients, and recordings collected in controlled environments. In this study, we enrolled 512 participants at two hospitals in Zambia, categorised into three groups: bacteriologically confirmed TB (TB+), symptomatic patients with other respiratory diseases (OR), and healthy controls (HC). Usable cough recordings with demographic and clinical data were obtained from 500 participants. Deep learning classifiers based on pre-trained speech foundation models were fine-tuned on cough recordings to predict diagnostic categories. The best-performing model, trained on 3-second audio clips, achieved an AUROC of 85.2% for distinguishing TB coughs from all other participants (TB+/ Rest) and 80.1% for TB+ versus symptomatic OR participants (TB+/ OR). Incorporating demographic and clinical features improved performance to 92.1% for TB+/ Rest and 84.2% for TB+/ OR. At a probability threshold of 0.38, the multimodal model reached 90.3% sensitivity and 73.1% specificity for TB+/ Rest, meeting WHO target product profile benchmarks for TB screening. Adversarial testing and stratified analyses shows that the model was robust to confounding factors including background noise, recording time, and device variability. These results demonstrate the feasibility of cough-based AI for TB screening in real-world, low-resource settings.
{IEEEkeywords}
Tuberculosis, TB, Acoustic biomarkers, Cough sound analysis, Artificial intelligence, Screening tools, Low-resource settings
1 Introduction
@@
Despite being preventable and curable, tuberculosis (TB) remains a major global health challenge, with 10.6 million new cases and 1.3 million deaths reported in 2023 [1]. A significant proportion of TB-related deaths, nearly one quarter of all estimated TB cases, stems from undiagnosed and untreated cases [1]. Systematic screening is essential to facilitate earlier diagnosis, reduce community transmission, and improve treatment outcomes [2], but current World Health Organisation (WHO)-approved screening tools are limited. Symptom-based screening offers only moderate sensitivity (71%) and specificity (64%) [3], while chest X-ray (CXR) requires costly infrastructure, with limited availability in many low-resource settings. C-reactive protein (CRP) testing shows diagnostic utility primarily among people living with HIV (PLHIV). Molecular WHO-recommended rapid diagnostic (mWRD), although valuable, are constrained by moderate sensitivity (69%), which is only recommended for adolescents and adults living with HIV and requires a robust sample transportation network [4, 5, 6].
@@
Advances in artificial intelligence (AI) offer opportunities to improve TB screening and diagnosis by enhancing accuracy, efficiency, and accessibility across various healthcare domains, including medical imaging [7, 8, 9], disease detection [10, 11, 12, 13], and sleep-disordered breathing monitoring [14]. AI-based CXR interpretation tools were recommended by WHO in 2021 as alternatives to human readers for TB screening [2], but their use remains constrained by equipment availability [5, 6]. These challenges highlight the need for alternative AI applications in TB screening [2], particularly those that exploit other diagnostic substrates such as audio signals and wearable sensor data, which may be more accessible and scalable in low-resource settings.
@@
An emerging alternative for TB screening is AI-based analysis of cough sounds. Cough is a hallmark symptom of TB, caused by inflammation of the airways, and is thought to contain disease-specific acoustic features distinguishable from coughs of other causes [15]. Clinicians already use cough characteristics such as tone, pitch, intensity, and duration of coughs, alongside associated symptoms, to guide diagnosis [16, 17]. AI-enabled cough analysis could formalise this approach, reduce variability, and scale to low-resource settings [18]. Recent studies have explored AI-enabled cough sound analysis as a screening tool for TB [15, 19, 20, 21]. However, prior research in this area has been limited by several weaknesses, including small sample sizes [15, 19], data biases due to inadequate inclusion of symptomatic individuals without TB [15, 19, 20], the use of relatively simple machine learning models [15, 22], and data collected in controlled, quiet environments that do not reflect real-world conditions [15, 21].
@@
In this study, we propose a novel AI-enabled cough sound analysis model specifically designed to overcome these limitations. We hypothesise that, with a sufficiently large and well-balanced dataset, machine learning models could learn to recognise spectral and temporal features in cough sounds that reliably distinguish TB from non-TB cases. Our contributions are threefold: (1) we present one of the largest and most balanced datasets of cough sounds for TB screening collected in noisy, real-world environments; (2) we demonstrate the effectiveness of speech foundation models for this task, achieving performance that meets WHO target product profile benchmarks; and (3) we systematically analyse robustness across devices, subgroups, and confounding factors, highlighting pathways for practical deployment in resource-limited settings. Our objective is to assess model performance in a high-burden setting, thereby advancing the development of practical and scalable tools for TB screening. Such tools could be particularly valuable in resource-limited health systems and in high-throughput contexts such as immigration screening centres.
2 Methods
2.1 Study design and participants
@@
This cross-sectional study follows the Standards for Reporting of Diagnostic Accuracy Studies (STARD) guidelines [23]. Ethical approval was obtained from the University of Zambia Biomedical Research Ethics Committee (approval number 3648-2023). Written informed consent was obtained from all participants prior to enrolment.
@@ 2.3 Model development
A multi-stage classification pipeline (Fig. 1) was developed using state-of-the-art pre-trained speech foundation models. These models, including Wav2Vec2 [25], WavLM [26], HuBERT [27], Data2Vec [28], and Whisper [29], accept raw audio waveform as input and generate contextualised audio embeddings that capture complex acoustic patterns. To adapt each model for three-way classification (TB+, OR, HC) in this study, we appended a classifier head comprising a global average pooling layer, dropout (rate = 0.5), and a softmax output layer. Fine-tuning was conducted using stratified 10-fold cross-validation with the following parameters: 5 training epochs, batch size = 8, learning rate = 3e-5, warm-up ratio = 0.1, and gradient accumulation over 8 steps.
@@ 2.5 Evaluation metrics
Table 2: Comparison of AUROC with different classifier architectures and audio input durations. Best results are in bold.
@@
3 Results
[…]
3.1 Different classifier architectures and audio durations
@@
TB+/ Rest (HIV+) Audio Alone 81.5% 75.4% 69.7% 73.2% 72.1% 74.1%
Audio + All Info 91.8% 81.9% 90.9% 94.7% 71.5% 87.8%
3.2 Incorporating demographic and clinical features
Using the Wav2Vec2-based classifier and 3-second audio input, we evaluated classification performance when demographic or clinical features (gender, age, BMI, or symptom presence) are available (Table 3). The ROC curves in different tasks are shown in Fig. 2. With audio alone, classification performance was consistently better for the TB+/ HC task than for TB+/ OR. Specificity was similar across both tasks (73.1% vs. 73.2%), but sensitivity (92.6% vs. 72.0%), PPV (71.8% vs. 66.7%), NPV (93.1% vs. 77.8%) and F1-score (80.9% vs. 69.1%) were all notably higher for TB+/ HC. These differences suggest greater acoustic overlap between TB and other symptomatic respiratory conditions than between TB and healthy controls. When the OR and HC groups are combined in the TB+/ Rest task, sensitivity is 82.3%, specificity is 73.2%, PPV is 82.0%, NPV is 73.5% and F1-score is 82.1%. Adding individual demographic or clinical features to audio inputs resulted in modest improvements across all tasks. The most substantial gains were observed when all features were combined. For TB+/ Rest, the inclusion of all supplemental information increased AUROC from 85.2% to 92.1%, sensitivity from 82.3% to 85.6%, and specificity from 73.2% to 83.1%. Similarly, F1-score improved from 82.1% to 86.9%. In the more challenging TB+/ OR task, where both groups were symptomatic, AUROC increased from 80.1% to 84.2%, and specificity from 73.2% to 83.1%, although sensitivity declined slightly from 72.0% to 71.3%. Overall, combining audio with demographic and clinical data led to consistent performance improvements across most metrics, with the strongest effects seen in the TB+/ Rest task.
@@ 3.3 HIV co-infection
3.4 Different classification thresholds
Table 4 shows model performance with different classification thresholds ranging from 0.36 to 0.50. For the TB+/ Rest task, a threshold of 0.38 offered a strong balance of sensitivity (90.3%) and specificity (73.1%), meeting the WHO Target Product Profile for TB screen tools, which recommended a minimum of >90% sensitivity and >70% specificity for screening tools in 2021 [31] and a minimum of >90% sensitivity and >60% specificity in 2025 [32]. Higher thresholds improved specificity (up to 83.1% at 0.50) at the cost of reduced sensitivity (85.6% at 0.50). For TB+/ OR, performance was generally lower than TB+/ Rest. At the 0.38 threshold, sensitivity was 80.6% and specificity was 73.1%. As the threshold increased, specificity improved (up to 83.1%) while sensitivity dropped (71.3% at 0.50). In the TB+/ HC task, the model achieved perfect sensitivity (100%) across all thresholds tested, with specificity increasing from 70.2% to 83.1% between thresholds 0.36 and 0.50.
Table 4: Comparison of results for the proposed system (audio + all demographic information) with different threshold scores. Bracketed figures show 95% confidence intervals based on 10,000 repeats. The results meeting the WHO target profile for TB screening tests are in bold.
@@ 3.4 Different classification thresholds
3.5 Potential confounders
Table 5: Effect of different noise training and testing conditions using the proposed audio classifier.
@@ 3.5 Potential confounders
Background Background 69.9%
Cough Sound Background 58.6%
@@
When both trained and tested using white noise as audio input, the model performed at around the chance level (AUROC 56.2%), indicating no meaningful discriminative capacity in the absence of cough sound (Table 5). In contrast, training and testing on non-cough acoustic background led to a significant increase in performance (AUROC = 69.9%), which suggests the presence of some background noise patterns associated with different participant groups. However, this performance remained well below that observed when trained and tested on cough sounds (AUROC = 85.2%), confirming that cough acoustics carry the primary discriminatory information. Notably, when the model was trained on cough sounds and tested on background noise, AUROC values again approached chance (58.6%), further underscoring the specificity of the learned representations to cough content rather than background artefacts.
@@
The spectral energy distribution of cough segments and background noise across groups can be visualised in the long-term average spectrum (LTAS) (Fig. 3). Cough LTAS profiles (left panel) were broadly similar across TB+, OR, and HC groups, with only minor differences: TB+ exhibiting slightly lower energy around 200 Hz and higher energy above 1.5 kHz. This indicates that group-level differences in cough acoustics are not prominently expressed in the long-term spectral domain, suggesting that temporal features may offer greater diagnostic value. In contrast, background noise LTAS (right panel) showed clearer group differences, with TB+ recordings exhibiting a flatter, higher-energy spectrum across frequencies, likely due to contextual (e.g., recording environment or timing) rather than physiological factors. However, adversarial testing (Table 5) showed that any such confounding is minimal, as near-chance performance was reported when models were trained on cough sounds and tested on background noise.
3.6 Mobile recordings
@@
Table 6: Performance of classifiers (with and without additional demographic information) using mobile phone recordings.
@@
4 Discussion
This study demonstrates the potential of AI-enabled cough sound analysis using speech foundation models to identify TB from short cough recordings. Our multimodal model achieved performance that meets the WHO target product profile benchmarks for TB screening tools [31, 32] in distinguishing TB-positive cases from individuals without TB (TB+/Rest, AUROC 92.1%, sensitivity 90.3%, specificity 73.1%). Based on a well-balanced cohort of 500 participants (201 TB+, 150 symptomatic non-TB, 149 healthy controls), these findings highlight the promise of cough-based AI as a scalable, non-invasive, and low-cost screening tool in high-burden settings with limited diagnostic infrastructure.
Compared with previous studies, our model achieved robust and reproducible performance despite real-world recording conditions. Direct comparisons are limited due to key methodological differences across studies. Prior research in this area has a number of weaknesses, including smaller or imbalanced datasets [15, 19], limited inclusion of symptomatic non-TB patients [15, 19, 20], the use of relatively simple machine learning models [15, 22], and data collection in controlled, quiet environments [15, 21], leading to potential bias and inflated accuracy. Our findings align with real-world clinical practice and prior studies [33], demonstrating that the inclusion of demographic and clinical parameters, such as age, gender, BMI, and symptom presence, can enhance model performance. For the TB+/ Rest task, model sensitivity improved substantially, primarily driven by the inclusion of asymptomatic healthy controls. As this subgroup lacks symptoms, the distinction between TB and non-TB cases becomes clearer, enabling the model to more accurately detect TB-positive individuals. In contrast, for the TB+/ OR task, where both groups were symptomatic, clinical features contributed primarily to improved specificity, allowing the model to better distinguish TB from other respiratory illnesses.
The lower performance of the proposed model in the TB+/ OR task compared to TB+/ HC highlights the clinical complexity of differentiating TB from other respiratory conditions. Including symptomatic non-TB participants is crucial for realistic rigorous performance assessment. Prior studies have often overlooked this group, or employed a small imbalanced dataset, which could lead to inflated accuracy estimates [15, 19, 20]. For example, Sharma et al. [21] reported highly variable classifier performance (AUROC 61–86%), largely due to training on imbalanced datasets (103 TB patients and 46 non-TB participants).